OpenStudy (anonymous):

Solve the equation by applying the squaring property of equality. Be sure to check all solutions in the original equation. ( If there is no solution, enter NO SOLUTION.)

OpenStudy (anonymous):

\[\sqrt{y-2}=y-4\]

OpenStudy (anonymous):

\[\sqrt{y-2}=y-4\] \[(\sqrt{y-2})^{2}=(y-4)^{2}\] \[y-2=y^{2}-8y+16\] \[y^{2}-9y+18=0\] And the last part I leave it to you. Just use the quadratic formula. : )

OpenStudy (anonymous):

Ok, thank you...don't leave , i might need you lol

OpenStudy (anonymous):

I got 6 & 3

OpenStudy (anonymous):

Yes, you are correct

OpenStudy (anonymous):

Yaaaaaay ! thank you, help me with the next one?

OpenStudy (anonymous):

OK, shoot it